OpenStudy (anonymous):

Use the squared identities to simplify 2sin2x cos2x. Check below.

OpenStudy (anonymous):

A. (1 - cos(4x))/4 B. (1 + 4cos(2x) + cos(4x))/4 C. (1 + cos(4x))/4 D. (1 + 4cos(2x) - cos(4x))/4

OpenStudy (anonymous):

I still can't figure it out. Can someone walk me through it?

OpenStudy (anonymous):

I'm only guessing B.

OpenStudy (anonymous):

2sin2xcos2x=sin4x

OpenStudy (anonymous):

Okay, I figured out that part... But how do I apply it do the question?

OpenStudy (anonymous):

I'm thinking A or C now!

OpenStudy (psymon):

C x_x Gah, I had to work a bit on that one. I'll work it out for ya, though, haha.

OpenStudy (anonymous):

A right?

OpenStudy (psymon):

\[2\sin(2x)\cos(2x) = \sin(4x)\] \[\sin(4x) = \sin(2x + 2x) \] \[\sin(2x+2x) = \sin(2x)\cos(2x) + \sin(2x)\cos(2x)\] \[2*(\frac{ 1-\cos(2x) }{ 2 })(\frac{ 1+\cos(2x) }{ 2 })\] \[2*(\frac{ 1-\cos ^{2}(2x) }{ 4 }) \] \[2*(\frac{ 1-\frac{ 1+\cos(4x) }{ 2 } }{ 4 })\] \[2(\frac{ \frac{ 2-1+\cos(4x) }{ 2} }{ 4 })\] \[2*(\frac{ 1+\cos(4x) }{ 8 }) = \frac{ 1+\cos(4x) }{ 4 }\]

OpenStudy (psymon):

*kicks trigonometry in the face*

OpenStudy (anonymous):

http://prntscr.com/48wqi4

OpenStudy (anonymous):

1-cos(4x) _________ 4